# Sketch2Sound: Controllable Audio Generation via Time-Varying Signals and Sonic Imitations

### v1

@@ I Introduction
The human voice is a gestural sonic instrument [12]: it allows us to realize sounds without having to perform any symbolic abstraction (i.e., putting a sound into words) beforehand. When humans communicate audio concepts to other people (rather than software), they typically combine descriptive language and vocal imitation [13, 14, 15]. In doing so, one approximates the audio by mapping the pitch, timbre, and temporal properties of the sound to those of the voice. This is a more natural method than describing the evolution of pitch, timing, and timbre via pure text descriptions [13].
@@
Our experiments show that Sketch2Sound can generate sounds that closely follow the input control signals (loudness, spectral centroid, and pitch/periodicity) from a vocal imitation while still achieving a high degree of adherence to a text prompt and an audio quality comparable to the text-only pre-trained model. We show that our median filtering technique leads to improved audio quality and text adherence when generating sounds from vocal imitations. We also show that, during inference, a user can arbitrarily specify a degree of temporal detail by choosing a median filter size, allowing them to navigate the trade-off between strict adherence to the vocal imitations and audio quality and text adherence.
@@
control signals text-only median (sz 10) 0.273 13.41 10.34 13.91 2.96 2.57 0.27
@@ II Method
We propose a method for conditioning an audio latent diffusion model on a set of interpretable, time-varying control signals that are suitable for multiple tasks. These include: generating variations of sounds, modifying existing sounds, and generating new sounds expressively via (optionally text-prompted) sonic imitations.
@@ II-A Time-varying control signals for sound objects
We choose the following three control signals to be used as conditioning for Sketch2Sound:
@@
  • Spectral Centroid is defined as the center of mass of the frequency spectrum for a given audio frame. Frames with a higher spectral centroid will be perceived as having a brighter timbre. To preprocess the centroid, we convert the signal from linear frequency space (i.e., Hz) to a continuous MIDI-like representation, scaled to roughly a $(0,1)$ range by dividing the input signal by $127$ (note G9, roughly 12.5kHz), which we found to stabilize the first steps of training.
@@ II-B Conditioning a latent audio DiT on time-varying control signals
We use a large pre-trained text-to-sound latent diffusion transformer (DiT), similar to the one described in [2, 21] (text-conditioned only, no timing conditioning) and adapt it to generate sounds conditioned on the time-varying control signals mentioned above. The latent diffusion model for text-to-sound generation has two parts: first, a variational autoencoder (VAE) compresses 48kHz mono audio to a sequence of continuous vectors of dim 64 at a rate of 40Hz. Then, a transformer decoder-only model is trained to generate new sequences of latents, which can be decoded into audio using the VAE decoder. This text-to-audio DiT was pre-trained on a large mix of proprietary, licensed sound effect datasets and publicly available CC-licensed general audio datasets. Once the model is pre-trained, we fine-tune it for 40k steps and adapt it to handle our time-varying control signals as conditioning.
Because the aforementioned time-varying control signals can be easily and efficiently extracted from any audio signal on the fly, we can fine-tune the pre-trained text-to-audio model in a self-supervised manner. Given any input audio signal, we extract the aforementioned three control signals (loudness, centroid, pitch) from the audio signal and use them as conditioning for the model during fine-tuning. The model is then fine-tuned using the same recipe used during training: learning the reverse diffusion process from a set of noisy latents with text conditioning, along with our proposed control conditioning.

### v2

@@ I Introduction
The human voice is a gestural sonic instrument [12]: it allows us to realize sounds without having to perform any symbolic abstraction (i.e., putting a sound into words) beforehand. When humans communicate audio concepts to other people, they typically combine language and vocal imitation [13, 14, 15], and recent work has shown its utility for query-by-example search of audio databases [16, 17, 18]. This is a more natural method than describing the evolution of pitch, timing, and timbre via pure text descriptions [13], and voice-driven sound synthesis interactions have been of interest long before modern generative modelling for their embodied capabilities[19, 20, 21, 22, 23, 24].
@@
Our experiments show that Sketch2Sound can generate sounds that closely follow the input control signals (loudness, spectral centroid, and pitch/periodicity) from a vocal imitation while still achieving a high degree of adherence to a text prompt and an audio quality comparable to the text-only pre-trained model. We show that our median filtering technique leads to improved audio quality and text adherence when generating sounds from vocal imitations. We also show that, during inference, a user can arbitrarily specify a degree of temporal detail by choosing a median filter size, allowing them to navigate the trade-off between strict adherence to the vocal imitations and audio quality + text adherence.
@@
control signals text-only median (sz 10) 0̱.273 13.41 10.34 13.91 2.96 2.57 0.27
@@ II Method
We propose a method for conditioning an audio latent diffusion model on a set of interpretable, time-varying control signals that are suitable tasks creating variations of sounds and generating new sounds expressively via text-prompted sonic imitations.
@@ II-A Time-varying control signals for sound objects
We use the following control signals as conditioning for Sketch2Sound:
@@
  • Spectral Centroid is defined as the center of mass of the frequency spectrum for a given audio frame. Frames with a higher centroid will be perceived as having a brighter timbre. To preprocess the centroid, we convert the signal from linear frequency space (i.e., Hz) to a continuous MIDI-like representation, scaled to roughly a $(0,1)$ range by dividing the input signal by $127$ (note G9, roughly 12.5kHz), which we found to stabilize the first steps of training.
@@ II-B Conditioning a latent audio DiT on time-varying control signals
Refer to Figure 1 for a visual overview of our approach. We use a large pre-trained text-to-sound latent diffusion transformer (DiT), similar to the one described in [2, 30] (text-conditioned only, no timing conditioning) and adapt it to generate sounds conditioned on the time-varying control signals mentioned above. The latent diffusion model for text-to-sound generation has two parts: first, a variational autoencoder (VAE) compresses 48kHz mono audio to a sequence of continuous vectors of dim 64 at a rate of 40Hz. Then, a transformer model is trained to generate new sequences of latents, which can be decoded into audio using the VAE decoder. This text-to-audio DiT was pre-trained on a large mix of proprietary, licensed sound effect datasets and publicly available CC-licensed general audio datasets. Once the model is pre-trained, we fine-tune it for 40k steps and adapt it to handle our time-varying control signals.
Because the time-varying control signals can be easily and efficiently extracted from any audio signal on the fly, we can fine-tune the pre-trained text-to-audio model in a self-supervised manner: Given any input audio signal, we extract the three control signals (loudness, centroid, pitch) from the audio signal and use them as conditioning for the model during fine-tuning. The model is then fine-tuned using the same recipe used during training: learning the reverse diffusion process from a set of noisy latents with text conditioning, along with our proposed control conditioning.